Vicky is now 4 times as old as Cristy. five years ago,the sum of their ages was 45. how old are they now
Answer by piggy278(1)

Vicky-->V
Cristy-->(V/4)
Five Years Ago
Vicky-->V-5
Cristy-->(V/4)-5
V-5+(V/4)-5=45
V-5+(V/4)=50
V+(V/4)=55
V=(V/4)x4
(V/4)x5=V+(V/4)
(V/4)=55/5=11#